OpenStudy (anonymous):

What is the tenth term of the geometric series if the first, second and third terms are 0.5, 1.0, 2.0, ... ?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[r=a_{2}/a_{1}=1/0.5=2\]

OpenStudy (anonymous):

You are starting at 0.5, and each time, you double what you had before. Just keep doubling until you get to the 10th term.

OpenStudy (anonymous):

the series can be written as An = 2*(An-1) or .5*2^n-1 so A10 = .5*2^9 = 256

OpenStudy (anonymous):

\[a_{10}=a _{1}*r^{9}=0.5*2^9=2^8=256\]

OpenStudy (anonymous):

its result is 512

OpenStudy (anonymous):

0.5 time 2^10= 2^9 = 512? 0r how?

OpenStudy (anonymous):

Result should be 512

OpenStudy (anonymous):

Oh, whoops. Should be 256. It's only multiplied 9 times. 0.5*2^9

OpenStudy (anonymous):

its \[0.5*2^{n-1} = 0.5*2^{10-1} = 0.5*2^{9} = 256\]

OpenStudy (anonymous):

Formula \[a_{n}=a _{1}*r^{n-1}\]

OpenStudy (anonymous):

but its result is 512